ABCD is a square with AB = 12. Point P is in an interior point whose distances to A, B, and to the side CD are equal. Find that distance.

Answer by checkley75(3666)

The triangle APB is an isosceles triangle with 45,45,90 degree angles.
The 2 equal sides are:
x^2+x^2=12^2
2x^2=144
x^2=144/2
x^2=72
x=sqrt72
x=8.4853 ans.